Charges-q, Q, -q are placed at  equal distance on a straight line, if the total potential energy of the system of 3 charges is 0,then find ratio Q/q
 
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
let the distance is d
then
2K*(-q)*Q/d + K*(-q)*(-q)/2d = 0
-2Q + q/2 =0
Q/q = ¼
Hope it clears........................................